A seed company planted a floral mosaic of a national flag. The perimeter of the flag is 460 feet. Determine the​ flag's length and width if the length is 70 feet greater than the width.
Answer by Boreal(15235)

Perimeter is 2L+2W
Therefore, half the perimeter is L+W
Half of 460 is 230
L+W=230, since they are half the perimeter
W+70+W=230, since L=70 more than W, or W+70
2W+70=230
2W=160, subtracting 70 from both sides
W=80 feet, dividing both sides of the equation by 2.
L=150 feet
2W=160 feet
2L=300 feet
Total is 460 feet check.